1.what is the expected number of days per month this year to two decimal places?
2.what if you exclude the months that end in " from the previous question?

Answer by oberobic(2304)

The expected number of days per month = sum of days per month / 12 months.
In 2011, there are 365 days.
365/12 = 30.42 days per month
.
If you exclude months that end in "R" that means you exclude 4 months: September, October, November, and December. They total 30+31+30+31 = 122 days.
.
365 - 122 = 243 days
.
243/8 = 30.38